OpenStudy (raffle_snaffle):

Given the points A(1,0,-1), B(2,-3,5), and C(0,5,1) in R^3 answer the following: i Find the angle, a (alfa), between the line segment AB and AC. ii Find the area of the triangle in space determined by A,B and C. iii Find parametric equation of the line L that contains the point A and is parallel to the line with symmetric equation x-3/4=y/2=z+2/-1

OpenStudy (raffle_snaffle):

i need to find the magnitude of the two lines then use tan(theta)?

OpenStudy (raffle_snaffle):

for part i

OpenStudy (dumbcow):

first use distance formula to get side lengths AB = sqrt46 AC = sqrt30 BC = 2sqrt21  i would use law of cosines to find angle alpha

OpenStudy (raffle_snaffle):

Do you have the formula for law of cosine

OpenStudy (raffle_snaffle):

distance formula is x^2+y^2+z^2?

OpenStudy (dumbcow):

\[c^{2} = a^{2} +b^{2}-2ab \cos C\] a,b,c are sides .... C is angle opp of side "c" distance formula \[d = \sqrt{(x_2 -x_1)^{2}+(y_2 -y_1)^{2}+(z_2 -z_1)^{2}}\]

OpenStudy (raffle_snaffle):

okay thanks

OpenStudy (zzr0ck3r):

you can use herons formula for the area

OpenStudy (raffle_snaffle):

herons formula?

OpenStudy (raffle_snaffle):

i am going to solve part i really quick

OpenStudy (zzr0ck3r):

k, ill post the formula

OpenStudy (raffle_snaffle):

thanks man

OpenStudy (zzr0ck3r):

\[\text{given the length of the three sides a, b and c }\\then\\Area = \sqrt{s(s-a)(s-b)(s-c)}\\where \space s \space = \frac{a+b+c}{2}\]

OpenStudy (raffle_snaffle):

The calc book they offer us at the college doesn't explain this material very well. I am trying to figure out which section in the book this question would be asked

OpenStudy (zzr0ck3r):

this is zach

OpenStudy (raffle_snaffle):

I know it's zach'

OpenStudy (zzr0ck3r):

ahh no need to tell me about the book then:P

OpenStudy (zzr0ck3r):

ah I c what your saying, this is calc 4 stuff

OpenStudy (zzr0ck3r):

So this question is not from the book?